OpenStudy (anonymous):

With the following equations, make a tables of values. Identify the x and y intercepts. Equation: y-3=0.5x

OpenStudy (anonymous):

Is your whole class here; I just answered this question 10 minutes ago! :)

OpenStudy (anonymous):

Its just me, but what are some other values for the table?

OpenStudy (anonymous):

Since the x is multiplied by 0.5, if you pick x-values that are multiples of 2, it will be easier to solve. Also, change the equation to y=0.5x+3. Then, when x=2, y=0.5(2)+3=1+3=4 x=4, y=0.5(4)+3=2+3=5 x=6, y=0.5(6)+3=3+3=6 etc.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You could also throw in a few negative values of x: x=-2, then y=0.5(-2)+3=-1+3=2 x=-4, then y=0.5(-4)+3=-2+3=1 etc Good?

OpenStudy (anonymous):

Yeah, thanks again.